One bar of tin alloy is 25% pure tin and another bar is 10% pure tin. How many pounds of each alloy must be used to make 75 pounds of a new alloy which is 20% pure tin?
Answer by fractalier(6550)

Call x the amount of 25% alloy used.
Then 75-x is the amount of 10% used.
The set up looks like this:
.25x + .10(75-x) = .20(75)
.25x + 7.5 - .10x = 15
.15x = 7.5
x = 50 pounds of 25%
75-x = 25 pounds of 10%